Chapter 22, Problem 050

At some instant the velocity components of an electron moving between two charged parallel plates are vx=1.9×105 m/s and vy=3.8×103 m/s. Suppose the electric field between the plates is uniform and given by E→=(120N/C)j︷. In unit-vector notation, what are (a) the electron’s acceleration in that field and (b) the electron’s velocity when its x coordinate has changed by 2.2 cm?
Answer #1

here,

vx = 1.9 * 10^5 m/s

vy = 3.8 * 10^3 m/s

the magnitude of feild , E = 120 N/C

a)

the electron's accelration , a = q * E /m

a = 1.6 * 10^-19 * 120 /( 9.1 * 10^-31) m/s^2

a = 2.11 * 10^13 m/s^2

b)

x = 2.2 cm = 0.022 m

the time taken travel in x-direction , t = x/vx

t = 0.022 /( 1.9 * 10^5) = 1.16 * 10^-7 s

the electron's velocity , vy' = vy + a * t

vy' = 1.9 * 10^5 + 2.11 * 10^13 * 1.16 * 10^-7 m/s

vy' = 2.64 * 10^6 m/s